# High-energy gamma-ray emission from memory-burdened primordial black holes

### v1

@@ Abstract
Theoretical studies on the memory-burden effect suggest that Primordial Black Holes (PBHs) with masses smaller than $10^{15}$ grams may be viable dark matter candidates and, consequently, be potential sources of high-energy particles in the present Universe. In this paper, we investigate the evaporation of memory-burdened PBHs into high-energy gamma-rays. Differently from previous analyses, we account for the attenuation of gamma-rays caused by their interaction with background radiation at energies above $10^{5}~{}{\rm GeV}$, as well as the secondary emission from the electromagnetic cascades generated during the propagation through extragalactic space. Performing a likelihood analysis with current gamma-ray data, we place new constraints on the parameter space of memory-burdened PBHs. Our results show that ultra-high-energy diffuse gamma-ray observations set more restrictive bounds than high-energy neutrino data, particularly in scenarios with a strong memory-burden suppression of the PBH evaporation.
@@ I Introduction
The semi-classical description of black hole evaporation assumes that the black hole remains classical throughout its lifetime Hawking (1975). This model may lack self-consistency, as notably known in the context of the information loss paradox Almheiri et al. (2021); Buoninfante et al. (2021). Specifically, Hawking’s original analysis neglects the influence of the emitted radiation on the quantum state of the black hole. However, this effect is expected to be significant when the energy of the emitted particles is comparable to the total energy of the black hole. Recently, Ref.s Dvali (2018); Dvali et al. (2020, 2024) have proposed that quantum back-reaction may significantly slow down black hole evaporation, as universally observed in quantum systems where the retained information resists decay. This effect, known as “memory burden”, would imply that PBHs with $M_{\rm PBH}\lesssim 10^{15}~{}{\rm g}$ could still be evaporating today, potentially leading to fascinating phenomenological consequences Alexandre et al. (2024); Thoss et al. (2024); Haque et al. (2024); Balaji et al. (2024); Barman et al. (2024b); Bhaumik et al. (2024); Barman et al. (2024c); Kohri et al. (2024); Jiang et al. (2024); Chianese et al. (2025); Zantedeschi and Visinelli (2024); Barker et al. (2024); Borah and Das (2025); Loc (2025); Basumatary et al. (2025); Athron et al. (2024); Bandyopadhyay et al. (2025); Calabrese et al. (2025); Boccia and Iocco (2025); Liu et al. (2025); Dvali et al. (2025); Montefalcone et al. (2025).
@@
In this paper, we thoroughly examine the emission of high-energy gamma-rays from the evaporation of memory-burdened PBHs, which may constitute a fraction of the DM component. Hence, we establish new limits on the parameter space of memory-burdened PBHs by analyzing the current upper bounds on the Ultra-High-Energy (UHE) diffuse gamma-ray flux from $10^{5}$ to $10^{11}$ GeV placed by several experimental collaborations, including CASA-MIA Chantell et al. (1997), KASCADE and KASCADE-Grande Apel et al. (2017), Pierre Auger Observatory Pie (2015); Aab et al. (2017), and Telescope Array Abbasi et al. (2019). Moreover, we revisit previous constraints Thoss et al. (2024); Haque et al. (2024) from Fermi-LAT Ackermann et al. (2015) and LHAASO Cao et al. (2023) measurements below $10^{6}$ GeV by taking into account the gamma-ray attenuation and the secondary emission from the electromagnetic cascades. At energies higher than $\sim 10^{5}~{}{\rm GeV}$, the interaction of gamma-rays with galactic and extragalactic background radiation becomes kinematically possible, resulting in the creation of electron-positron pairs. This process prevents a fraction of gamma-rays from reaching the Earth, thus attenuating the gamma-ray flux from evaporating PBHs. Furthermore, electrons and positrons undergo interactions such as Inverse Compton scattering with background photons, bremsstrahlung, and synchrotron radiation in the galactic and intergalactic magnetic fields. These interactions, which occur multiple times during propagation, generate electromagnetic cascades that eventually produce a secondary flux of low-energy photons. We demonstrate here that this component represents the dominant contribution in the Fermi-LAT energy range, as the primary gamma-ray emission is strongly suppressed by attenuation.
@@ II Memory-burdened primordial black holes
Figure 1: Diffuse gamma-ray flux. The different colors correspond to different combinations of the initial PBH mass $M_{\rm PBH}$ and the memory-burden parameter $k$, taking $f_{\rm PBH}=10^{-6}$. The solid/dashed lines show the prompt galactic gamma-ray flux (see Eq. (13)) with/without the effect of gamma-ray attenuation, while the dot-dashed lines show the universal spectrum of the secondary emission from extragalactic electromagnetic cascades (see Eq. (16)). The fluxes are averaged over $\Delta\Omega=4\pi$. From left to right, the data points correspond to the measurements of Fermi-LAT Ackermann et al. (2015) and LHAASO Cao et al. (2023) (inner and outer regions in gray and black colors, respectively).
@@
Here, $\mathcal{G}\approx 3.8$ accounts for gray-body effects due to gravitational back-scattering Page and Hawking (1976), while $g_{\rm SM}\approx 102.6$ represents the number of relativistic degrees of freedom at the temperature $T_{\rm H}$ in the Standard Model Mazde and Visinelli (2023). In the conventional scenario, this evaporation process continues until the entire PBH mass is transformed into radiation. The time required for complete evaporation results to be
@@
with $M_{\rm P}=(8\pi G)^{-1/2}$ being the reduced Planck mass. Consequently, the total evaporation time extends to
@@
For this analysis, we assume $q=1/2$ as the memory-burden effect is expected to become significant when half of the initial PBH mass has been lost. It is worth noting that the precise choice of $q$ only influences how our results are interpreted in terms of the original PBH mass. Moreover, we assume an instantaneous transition from the semi-classical regime to the memory-burdened one. Recent studies Dvali et al. (2025); Montefalcone et al. (2025) have demonstrated that a non-instantaneous transition could significantly alter the PBH evolution throughout the entire history of the Universe. However, as will be discussed in the following sections, the gamma-ray constraints are primarily based on the current galactic emission. Consequently, the limits we report in this paper can be rescaled to account for the specific suppression of PBH evaporation in different non-instantaneous models. A dedicated analysis will be explored in detail in future work.
@@ III The high-energy gamma-ray emission
We analyze the flux of high-energy gamma-rays produced by a population of memory-burdened PBHs, which contribute a fraction $f_{\rm PBH}=\Omega_{\rm PBH}/\Omega_{\rm DM}$ to the total DM density of the Universe. We consider a monochromatic PBH mass spectrum within the range $10^{-1}~{}{\rm g}\leq M_{\rm PBH}\leq 10^{9}~{}{\rm g}$.
@@
$\frac{{\rm d}^{2}N^{\rm mb}_{\gamma}}{{\rm d}E{\rm d}t}=S(M_{\rm PBH})^{-k}\,\frac{{\rm d}^{2}N_{\gamma}}{{\rm d}E{\rm d}t}\,.$ (12)
@@
with scale radius $r_{s}=25~{}{\rm kpc}$ and $\rho_{s}=0.23~{}{\rm GeV/cm^{3}}$ providing a local DM density of $0.4~{}{\rm GeV/cm^{3}}$ Iocco et al. (2015); Benito et al. (2019, 2021). This choice also ensures a consistent comparison with the constraints recently placed using high-energy neutrinos Chianese et al. (2025). The optical depth is computed following Ref.s Capanema et al. (2021); Chianese et al. (2021). We take into into account both the homogeneous cosmic microwave background (CMB) and the galactic starlight and infrared radiation as taken from the GALPROPv54 code Porter et al. (2017). Although the galactic radiation varies across the Milky Way, and peaks toward the galactic center and along the galactic plane, the angular dependence of the J-factor is primarily determined by the DM density profile. We have that, for $\Delta\Omega=4\pi$, the averaged J-factor at $10^{6}~{}{\rm GeV}$ is $\mathcal{J}=7.84\times 10^{21}~{}{\rm GeV/cm^{2}/sr}$ and $\mathcal{J}=2.22\times 10^{22}~{}{\rm GeV/cm^{2}/sr}$ with and without the gamma-ray attenuation, respectively.
@@
where the first term fixes the normalization with $\Omega_{\rm DM}=0.264$ and $\rho_{\rm c}=4.79\times 10^{-6}~{}{\rm GeV/cm^{3}}$ Aghanim et al. (2020), and the last term corresponds to an isotropic spectrum with a nearly universal shape, which depends on the CMB and the extragalactic background light (EBL) Berezinsky and Kalashev (2016). We numerically compute this spectrum by means of the $\gamma$-CascadeV4 code Blanco (2019); Capanema and Blanco (2025), providing as input the injected gamma-ray spectrum given in Eq. (12) and assuming the best-fit EBL model from Ref. Saldana-Lopez et al. (2021). We take into account the time evolution $M_{\rm PBH}(t)$ of the PBH mass and integrate up to a redshift $z_{\rm max}=10$, which is the maximum allowed value by the $\gamma$-CascadeV4 code.
In Fig. 1 we show the prompt galactic emission defined in Eq. (13) (solid lines) and the secondary extragalactic emission defined in Eq. (16) (dot-dashed lines) in case of three different choices for the initial PBH mass and the parameter $k$, assuming $f_{\rm PBH}=10^{-6}$. In order to highlight the effect of gamma-ray attenuation, which is relevant from $10^{5}$ to $10^{9}$ GeV, we also show the prompt galactic emission without attenuation with thin dashed lines. The data points correspond to the isotropic gamma-ray background (IGRB) measured by Fermi-LAT Ackermann et al. (2015) (red points below $10^{3}$ GeV) and to the diffuse emission from an inner region ($15^{\circ}<l<125^{\circ}$ and $|b|<5^{\circ}$, gray points) and an outer region ($125^{\circ}<l<235^{\circ}$ and $|b|<5^{\circ}$, black points) of the Milky Way recently measured by LHAASO Cao et al. (2023). The smaller the PBH mass, the higher the energies of the gamma-ray emitted by the galactic PBH distribution, which is typically well beyond the Fermi-LAT energy range. Hence, we find that the Fermi-LAT and LHAASO telescopes can mainly probe the secondary extragalactic emission and the prompt galactic emission, respectively.
Figure 2: Integrated gamma-ray flux. The different colors correspond to different combinations of the initial PBH mass $M_{\rm PBH}$ and the memory-burden parameter $k$, taking $f_{\rm PBH}=10^{-6}$. These fluxes are obtained from Eq. (17) with/without taking into account the gamma-ray attenuation (solid/dashed lines). The fluxes are averaged over $\Delta\Omega=4\pi$. Dray data points are the upper limits on the UHE gamma-ray flux placed by CASA-MIA Chantell et al. (1997), KASCADE and KASCADE-Grande Apel et al. (2017) at 90% CL, and by Pierre Auger Observatory (PAO) Pie (2015); Aab et al. (2017) and Telescope Array Surface Detector (TASD) Abbasi et al. (2019) at 95% CL.
@@
in case of the three benchmark scenarios previously discussed. In these energy range, the gamma-ray emission is dominated by the galactic component, while the extragalactic contribution is highly negligible. As before, the solid and dashed lines represent the flux with and without the gamma-ray attenuation, respectively. In the figure, we also report a collection of upper bounds placed by several experiments Chantell et al. (1997); Apel et al. (2017); Pie (2015); Aab et al. (2017); Abbasi et al. (2019).
@@ IV Results
Figure 3: Gamma-ray constraints on memory-burdened PBHs as viable DM candidates. Left: the colored lines refer to gamma-ray limits placed at 95% C.L. in the $M_{\rm PBH}$-$k$ plane with $f_{\rm PBH}=1$, in case of Fermi-LAT Ackermann et al. (2015) (solid blue line), LHAASO Cao et al. (2023) (dashed yellow line) and UHE diffuse gamma-ray Chantell et al. (1997); Apel et al. (2017); Pie (2015); Aab et al. (2017); Abbasi et al. (2019) (dot-dashed red line) data. The white area represents memory-burdened PBHs that can serve as viable DM candidates ($f_{\rm PBH}=1$), whereas the hatched region indicates PBHs that have fully evaporated in cosmological times. Previous gamma-ray Haque et al. (2024) and neutrino limits Chianese et al. (2025) are shown by the thin black lines with dotted and dashed styles, respectively. Right: ratio of the limits reported in the left panel to the gamma-ray constraint previously obtained in Ref. Haque et al. (2024), which is displayed as reference with the horizontal dotted line.
Figure 4: Gamma-ray constraints on the PBH abundance. Constraints at 95% CL in the $M_{\rm PBH}$-$f_{\rm PBH}$ plane for different values of the memory-burden parameter $k$ ($k=1.0$ left panel, $k=2.0$ middle panel, and $k=3.0$ right panel). The different shading lines refer to different data samples: Fermi-LAT Ackermann et al. (2015) (solid blue line), LHAASO Cao et al. (2023) (dashed yellow line), UHE diffuse gamma-ray Chantell et al. (1997); Apel et al. (2017); Pie (2015); Aab et al. (2017); Abbasi et al. (2019) (dot-dashed red line), and HE neutrinos Chianese et al. (2025). The hatched regions indicates PBHs that have fully evaporated in cosmological times.
We analyze the data reported by each experimental collaboration discussed in Fig. 1 and 2 to place constraints on the PBH abundance $f_{\rm PBH}$ as a function of $\vec{\theta}_{\rm PBH}=(M_{\rm PBH},k)$. For each experiment, we consider the following background-agnostic likelihood function
@@
We highlight a few key aspects of our results. First, we find that the constraints we place with Fermi-LAT and LHAASO data significantly differ from the ones obtained in previous analyses Thoss et al. (2024); Haque et al. (2024). For $k\lesssim 3.0$, they are weaker than the bounds placed by Ref. Haque et al. (2024), while for $k\gtrsim 3.0$ the Fermi-LAT limit results to be more stringent. This is due to the fact that the gamma-ray attenuation and the secondary emission from extragalactic electromagnetic cascades were not previously considered. We also note that the LHAASO telescope probes the scenario of memory-burden PBHs up to $k\simeq 4.4$. Indeed, higher values for the parameter $k$ correspond to a prompt galactic gamma-ray flux above $10^{6}~{}{\rm GeV}$, that is outside the LHAASO energy range. Second, the new limits imposed by the UHE gamma-ray data provide the most stringent constraints on the PBH parameter space, particularly for large values of the parameter $k$. Lastly, our results enable a consistent comparison between gamma-ray and neutrino constraints, as they are both derived under the same assumptions, e.g. for the DM distribution and the likelihood analysis. We robustly demonstrate that the current neutrino observations provide stronger constraints than gamma-ray data for $1.2\lesssim k\lesssim 1.5$ only.

### v2

@@ Abstract
Theoretical studies on the memory-burden effect suggest that Primordial Black Holes (PBHs) with masses smaller than $10^{15}$ grams may be viable dark matter candidates and, consequently, be potential sources of high-energy particles in the present Universe. In this paper, we investigate the evaporation of memory-burdened PBHs into high-energy gamma-rays. Differently from previous analyses, we account for the attenuation of gamma-rays caused by their interaction with background radiation at energies above $10^{5}~{\rm GeV}$, as well as the secondary emission from the electromagnetic cascades generated during their propagation through extragalactic space. Performing a likelihood analysis with current gamma-ray data, we place new constraints on the parameter space of memory-burdened PBHs. Our results show that ultra-high-energy diffuse gamma-ray observations set more restrictive bounds than high-energy neutrino data, particularly in scenarios with a strong memory-burden suppression of the PBH evaporation.
@@ I Introduction
The semi-classical description of black hole evaporation assumes that the black hole remains classical throughout its lifetime Hawking (1975). This model may lack self-consistency, as notably known in the context of the information loss paradox Almheiri et al. (2021); Buoninfante et al. (2021). Specifically, Hawking’s original analysis neglects the influence of the emitted radiation on the quantum state of the black hole. However, this effect is expected to be significant when the energy of the emitted particles is comparable to the total energy of the black hole. Recently, Ref.s Dvali (2018); Dvali et al. (2020, 2024) have proposed that quantum back-reaction may significantly slow down the black hole evaporation, as universally observed in quantum systems where the retained information resists decay. This effect, known as “memory burden”, would imply that PBHs with $M_{\rm PBH}\lesssim 10^{15}~{\rm g}$ could still be evaporating today, potentially leading to fascinating phenomenological consequences Alexandre et al. (2024); Thoss et al. (2024); Haque et al. (2024); Balaji et al. (2024); Barman et al. (2024b); Bhaumik et al. (2024); Barman et al. (2024c); Kohri et al. (2024); Jiang et al. (2024); Chianese et al. (2025); Zantedeschi and Visinelli (2024); Barker et al. (2024); Borah and Das (2025); Loc (2025); Basumatary et al. (2025); Athron et al. (2024); Bandyopadhyay et al. (2025); Calabrese et al. (2025); Boccia and Iocco (2025); Liu et al. (2025); Dvali et al. (2025); Montefalcone et al. (2025).
@@
In this paper, we thoroughly examine the diffuse emission of high-energy gamma-rays from the evaporation of memory-burdened PBHs, which may constitute a fraction of the DM component. Hence, we establish new limits on the parameter space of memory-burdened PBHs by analyzing the current upper bounds on the Ultra-High-Energy (UHE) diffuse gamma-ray flux from $10^{5}$ to $10^{11}$ GeV placed by several experimental collaborations, including CASA-MIA Chantell et al. (1997), KASCADE and KASCADE-Grande Apel et al. (2017), Pierre Auger Observatory Aab et al. (2015, 2017), and Telescope Array Abbasi et al. (2019). Moreover, we revisit previous constraints Thoss et al. (2024); Haque et al. (2024) from Fermi-LAT Ackermann et al. (2015) and LHAASO Cao et al. (2023) measurements below $10^{6}$ GeV by taking into account the gamma-ray attenuation and the secondary emission from the electromagnetic cascades. At energies higher than $\sim 10^{5}~{\rm GeV}$, the interaction of gamma-rays with galactic and extragalactic background radiation becomes kinematically possible, resulting in the creation of electron-positron pairs. This process prevents a fraction of gamma-rays from reaching the Earth, thus attenuating the gamma-ray flux from evaporating PBHs. Furthermore, electrons and positrons undergo interactions such as Inverse Compton scattering with background photons, bremsstrahlung, and synchrotron radiation in the galactic and intergalactic magnetic fields. These interactions, which occur multiple times during propagation, generate electromagnetic cascades that eventually produce a secondary flux of low-energy photons. We demonstrate here that this component represents the dominant contribution in the Fermi-LAT energy range, as the primary gamma-ray emission is strongly suppressed by attenuation.
@@ II Memory-burdened primordial black holes
Figure 1: Diffuse gamma-ray flux. The different colors correspond to different combinations of the initial PBH mass $M_{\rm PBH}$ and the memory-burden parameter $k$, taking $f_{\rm PBH}=10^{-6}$. The solid/dashed lines show the prompt galactic gamma-ray flux (see Eq. (13)) with/without the effect of gamma-ray attenuation, while the dot-dashed lines show the universal spectrum of the secondary emission from extragalactic electromagnetic cascades (see Eq. (16)). The fluxes are averaged over $\Delta\Omega=4\pi$. From left to right, the data points correspond to the isotropic gamma-ray background measured by Fermi-LAT Ackermann et al. (2015) and the galactic emission observed by LHAASO Cao et al. (2023) (inner and outer regions in gray and black colors, respectively).
@@
Here, $\mathcal{G}\approx 3.8$ accounts for gray-body effects due to gravitational back-scattering Page and Hawking (1976), while $g_{\rm SM}\approx 102.6$ represents the number of relativistic degrees of freedom at the temperature $T_{\rm H}$ in the Standard Model Mazde and Visinelli (2023). In the conventional scenario, the evaporation process continues until the entire PBH mass is transformed into radiation. The time required for complete evaporation results to be
@@
with $M_{\rm P}=1/\sqrt{8\pi G}$ being the reduced Planck mass. Consequently, the total evaporation time extends to
@@
For this analysis, we assume $q=1/2$ as the memory-burden effect is expected to become significant when half of the initial PBH mass has been lost. It is worth noting that the precise choice of $q$ only influences how our results are interpreted in terms of the original PBH mass. Moreover, we assume an instantaneous transition from the semi-classical regime to the memory-burdened one. Recent studies Dvali et al. (2025); Montefalcone et al. (2025) have demonstrated that a non-instantaneous transition could significantly alter the PBH evolution throughout the entire history of the Universe. However, as will be discussed in the following sections, the gamma-ray constraints are primarily based on the current galactic emission. Consequently, the limits we report in this paper can be rescaled to account for the specific suppression of PBH evaporation in different non-instantaneous models. A dedicated analysis will be provided in future work.
@@ III The high-energy gamma-ray emission
We analyze the flux of high-energy gamma-rays produced by a population of memory-burdened PBHs, which contribute a fraction $f_{\rm PBH}=\Omega_{\rm PBH}/\Omega_{\rm DM}$ to the total DM density of the today’s Universe. We consider a monochromatic PBH mass spectrum within the range $10^{-1}~{\rm g}\leq M_{\rm PBH}\leq 10^{9}~{\rm g}$.
@@
$\frac{{\rm d}^{2}N^{\rm mb}_{\gamma}}{{\rm d}E{\rm d}t}=\frac{1}{S(M_{\rm PBH})^{k}}\,\frac{{\rm d}^{2}N_{\gamma}}{{\rm d}E{\rm d}t}\,.$ (12)
@@
with scale radius $r_{s}=25~{\rm kpc}$ and $\rho_{s}=0.23~{\rm GeV/cm^{3}}$ providing a local DM density of $0.4~{\rm GeV/cm^{3}}$ Iocco et al. (2015); Benito et al. (2019, 2021). This choice also ensures a consistent comparison with the constraints recently placed using high-energy neutrinos Chianese et al. (2025). The optical depth in Eq. (14) is computed following Ref.s Capanema et al. (2021); Chianese et al. (2021). We take into into account both the homogeneous cosmic microwave background (CMB) and the galactic starlight and infrared radiation as taken from the GALPROPv54 code Porter et al. (2017). Although the galactic radiation varies across the Milky Way, and peaks toward the galactic center and along the galactic plane, the angular dependence of the J-factor is primarily determined by the DM density profile. We find that, for $\Delta\Omega=4\pi$, the averaged J-factor at $10^{6}~{\rm GeV}$ is $\mathcal{J}=7.84\times 10^{21}~{\rm GeV/cm^{2}}$ and $\mathcal{J}=2.22\times 10^{22}~{\rm GeV/cm^{2}}$ with and without the gamma-ray attenuation, respectively.
@@
where the first term fixes the normalization with $\Omega_{\rm DM}=0.264$ and $\rho_{\rm c}=4.79\times 10^{-6}~{\rm GeV/cm^{3}}$ Aghanim et al. (2020), and the last term corresponds to an isotropic spectrum with a nearly universal shape, which depends on the CMB and the extragalactic background light (EBL) Berezinsky and Kalashev (2016). We numerically compute this spectrum by means of the $\gamma$-CascadeV4 code Blanco (2019); Capanema and Blanco (2025), providing as input the injected gamma-ray spectrum given in Eq. (12) and assuming the best-fit EBL model from Ref. Saldana-Lopez et al. (2021). We take into account the time evolution $M_{\rm PBH}(t)$ of the PBH mass and integrate up to a redshift $z_{\rm max}=10$, which is the maximum allowed value by the $\gamma$-CascadeV4 code. By varying $z_{\rm max}$ from 1 to 10, we have verified that the secondary extragalactic emission remains unchanged at $E>10^{2}~{\rm GeV}$. This indicates that contributions from higher redshifts would have no impact on the Fermi-LAT limits which are mainly driven by the flux at these energies.
In Fig. 1 we show the prompt galactic emission defined in Eq. (13) (solid lines) and the secondary extragalactic emission defined in Eq. (16) (dot-dashed lines) in case of three different choices for the initial PBH mass and the parameter $k$, assuming $f_{\rm PBH}=10^{-6}$. In order to highlight the effect of gamma-ray attenuation, which is relevant from $10^{5}$ to $10^{9}$ GeV, we also show the prompt galactic emission without attenuation with thin dashed lines. The data points correspond to the isotropic gamma-ray background (IGRB) measured by Fermi-LAT Ackermann et al. (2015) (black points below $10^{3}$ GeV) and to the diffuse emission from an inner region ($15^{\circ}<l<125^{\circ}$ and $|b|<5^{\circ}$, gray points) and an outer region ($125^{\circ}<l<235^{\circ}$ and $|b|<5^{\circ}$, black points) of the Milky Way recently measured by LHAASO Cao et al. (2023). We find that the smaller the PBH mass, the higher the energies of the gamma-ray emitted by the galactic PBH distribution, which is typically well beyond the Fermi-LAT energy range. This implies that the Fermi-LAT and LHAASO telescopes can mainly probe the secondary extragalactic emission and the prompt galactic emission, respectively.
Figure 2: Integrated gamma-ray flux. The different colors correspond to different combinations of the initial PBH mass $M_{\rm PBH}$ and the memory-burden parameter $k$, taking $f_{\rm PBH}=10^{-6}$. These fluxes are obtained from Eq. (17) with/without taking into account the gamma-ray attenuation (solid/dashed lines). The fluxes are averaged over $\Delta\Omega=4\pi$. Gray data points are the upper limits on the UHE gamma-ray flux placed by CASA-MIA Chantell et al. (1997), KASCADE and KASCADE-Grande Apel et al. (2017) at 90% CL, and by Pierre Auger Observatory (PAO) Aab et al. (2015, 2017) and Telescope Array Surface Detector (TASD) Abbasi et al. (2019) at 95% CL.
@@
in case of the three benchmark scenarios previously discussed. In this energy range, the gamma-ray emission is dominated by the galactic component, while the extragalactic contribution is highly negligible. As before, the solid and dashed lines represent the flux with and without the gamma-ray attenuation, respectively. In the figure, we also report a collection of upper bounds placed by several experiments Chantell et al. (1997); Apel et al. (2017); Aab et al. (2015, 2017); Abbasi et al. (2019).
@@ IV Results
Figure 3: Gamma-ray constraints on memory-burdened PBHs as viable DM candidates. Left: the colored lines refer to gamma-ray limits placed at 95% C.L. in the $M_{\rm PBH}$-$k$ plane with $f_{\rm PBH}=1$, in case of Fermi-LAT Ackermann et al. (2015) (solid blue line), LHAASO Cao et al. (2023) (dashed yellow line) and UHE diffuse gamma-ray Chantell et al. (1997); Apel et al. (2017); Aab et al. (2015, 2017); Abbasi et al. (2019) (dot-dashed red line) data. The white area represents memory-burdened PBHs that can serve as viable DM candidates ($f_{\rm PBH}=1$), whereas the hatched region indicates PBHs that have fully evaporated in cosmological times. Previous gamma-ray Thoss et al. (2024) and neutrino limits Chianese et al. (2025) are shown by the thin black lines with solid and dashed styles, respectively. Right: ratio of the gamma-ray limits on $M_{\rm PBH}$ reported in the left panel to the neutrino constraint previously obtained in Ref. Chianese et al. (2025), which is displayed as reference with the horizontal dashed line.
Figure 4: Gamma-ray constraints on the PBH abundance. Constraints at 95% CL in the $M_{\rm PBH}$-$f_{\rm PBH}$ plane for different values of the memory-burden parameter $k$ ($k=1.0$ left panel, $k=2.0$ middle panel, and $k=3.0$ right panel). Lines with different shading refer to different data samples: Fermi-LAT Ackermann et al. (2015) (solid blue line), LHAASO Cao et al. (2023) (dashed yellow line), UHE diffuse gamma-ray Chantell et al. (1997); Apel et al. (2017); Aab et al. (2015, 2017); Abbasi et al. (2019) (dot-dashed red line), and high-energy neutrinos Chianese et al. (2025). The hatched regions indicates PBHs that have fully evaporated in cosmological times.
We analyze the data reported by each experimental collaboration reported in Fig.s 1 and 2 to place constraints on the PBH abundance $f_{\rm PBH}$ as a function of the parameters $\vec{\theta}_{\rm PBH}=(M_{\rm PBH},k)$. We assume that the data are well explained by the gamma-ray emission from astrophysical sources, while the PBH evaporation provides an additional contribution (see e.g. Ref.s Ajello et al. (2015); Ambrosone et al. (2021); Roth et al. (2021) and Lipari and Vernetto (2018); De La Torre Luque et al. (2025) for the astrophysical interpretation of the Fermi-LAT and LHAASO observations, respectively]). For each experiment, we therefore consider the following background-agnostic likelihood function
@@
We highlight a few key aspects of our results. First, we find that the constraints we place with Fermi-LAT and LHAASO data significantly differ from the ones obtained in previous analyses. For $k\lesssim 3.0$, they are weaker than the bounds placed by Ref. Thoss et al. (2024), while for $k\gtrsim 3.0$ the Fermi-LAT limit results to be more stringent. This is due to the fact that the gamma-ray attenuation and the secondary emission from extragalactic electromagnetic cascades were not previously considered. We also note that the LHAASO telescope probes the scenario of memory-burden PBHs up to $k\simeq 4.4$. Higher values for the parameter $k$ correspond to a prompt galactic gamma-ray flux above $10^{6}~{\rm GeV}$, which is outside the LHAASO energy range. Second, the new limits imposed by the UHE gamma-ray data provide the most stringent constraints on the PBH parameter space, particularly for large values of the parameter $k$. Lastly, our results enable a consistent comparison between gamma-ray and neutrino constraints, as they are both derived under the same assumptions, e.g. for the DM distribution and the likelihood analysis. We robustly demonstrate that the current neutrino observations provide stronger constraints than gamma-ray data for $1.2\lesssim k\lesssim 1.5$ only.